# Quantum Wave Simulation with Sources and Loss Functions

### v1

@@ Appendix A Estimating Loss Functions
Let $\mathbf{a},\mathbf{b}\in\mathbb{C}^{L}$. Given a projector $\mathbf{P}_{\mathcal{S}}=\text{diag}(p_{1},p_{2},\dots,p_{L})$, which projects onto a subspace $\mathcal{S}\subseteq\mathbb{C}^{L}$, where each $p_{i}\in\{0,1\}$ and $d=\sum_{i=1}^{L}p_{i}$, and let there be a state preparation oracle $\mathcal{U}$ that prepares the quantum state |$\ket{\phi}=[\mathbf{a};\mathbf{b}]/||[\mathbf{a};\mathbf{b}]||$ |; then, if either $d=\mathcal{O}(\text{polylog}(L))$ or $L-d=\mathcal{O}(\text{polylog}(L))$, there exists a quantum algorithm that efficiently estimates the $l_{2,\mathcal{S}}$-norm between $\mathbf{a}$ and $\mathbf{b}$ on $\mathcal{S}$ as
@@ A.1 Proof of Lemma 6
Here we consider the problem of estimating the $l_{2}$-norm between two vectors $\mathbf{a},\mathbf{b}\in\mathbb{C}^{L}$ on a subspace $\mathcal{S}\subseteq\mathbb{C}^{L}$ defined through the projector $\mathbf{P}_{\mathcal{S}}=\text{diag}(p_{1},p_{2},\dots,p_{L})$, where each $p_{i}\in\{0,1\}$ and $d=\sum_{i=1}^{L}p_{i}$. We will prove that the problem can be solved with probability at least $1-\delta$ within error $\epsilon$ by $\mathcal{O}(\log(1/\delta)/\epsilon)$ calls to a state preparation oracle $\mathcal{U}$ (along with its inverse and controlled versions), which prepares a quantum state |$\ket{\phi}=[\mathbf{a};\mathbf{b}]/||[\mathbf{a};\mathbf{b}]||$ | if either $d=\mathcal{O}(\text{polylog}(L))$ or $L-d=\mathcal{O}(\text{polylog}(L))$.
@@
By introducing |$\ket{\phi}=[\mathbf{a};\mathbf{b}]/||[\mathbf{a};\mathbf{b}]||$ |, we can write equation (37) as an expected value estimation
@@
To estimate $l_{2,\mathcal{S}}^{2}$ with optimal precision scaling beyond the central limit, we will introduce an auxiliary qubit to double the state space dimension. In this augmented space we can find a unitary decomposable observable $\hat{\mathbf{O}}$, whose expectation value with respect to the augmented state |$\ket{\psi}=[\mathbf{P}_{\mathcal{S}}\mathbf{a};\mathbf{P}_{\mathcal{S}}\mathbf{b};(\mathds{1}-\mathbf{P}_{\mathcal{S}})\mathbf{a};(\mathds{1}-\mathbf{P}_{\mathcal{S}})\mathbf{b}]/||[\mathbf{a};\mathbf{b}]||$ | is equal to $l_{2,\mathcal{S}}$, such that
$l_{2,\mathcal{S}}^{2}=\braket{\psi}{\hat{\mathbf{O}}}{\psi}||[\mathbf{a};\mathbf{b}]||^{2}.$ (40)
We first demonstrate that there exists a unitary initialization procedure for |$\ket{\psi}$ given an initialization oracle for |$\ket{\phi}$ and discuss the efficiency of its implementation. We then explicitly construct $\hat{\mathbf{O}}$, which can be decomposed into $4$ commuting unitary observables independent of $L$ and $\mathcal{S}$. We conclude by analyzing the overall complexity of our algorithm.
@@
where $n=log_{2}L$. We further define the unitary matrix $\hat{\mathbf{P}}_{\mathcal{S}}\in\mathbb{C}^{4L\times 4L}$ constructed from $\mathbf{P}_{\mathcal{S}}$ and its complement subspace $\mathds{1}-\mathbf{P}_{\mathcal{S}}$ as
$\hat{\mathbf{P}}_{\mathcal{S}}=\begin{bmatrix}\mathbf{P}_{\mathcal{S}}&0&(\mathds{1}-\mathbf{P}_{\mathcal{S}})&0\\ 0&\mathbf{P}_{\mathcal{S}}&0&(\mathds{1}-\mathbf{P}_{\mathcal{S}})\\ (\mathds{1}-\mathbf{P}_{\mathcal{S}})&0&\mathbf{P}_{\mathcal{S}}&0\\ 0&(\mathds{1}-\mathbf{P}_{\mathcal{S}})&0&\mathbf{P}_{\mathcal{S}}\end{bmatrix}.$ (42)
@@
|$\ket{\psi}=\hat{\mathbf{P}}_{\mathcal{S}}(I\otimes\mathbf{U})\ket{0^{n+1}}=[\mathbf{P}_{\mathcal{S}}\mathbf{a};\mathbf{P}_{\mathcal{S}}\mathbf{b};(\mathds{1}-\mathbf{P}_{\mathcal{S}})\mathbf{a};(\mathds{1}-\mathbf{P}_{\mathcal{S}})\mathbf{b}]/||[\mathbf{a};\mathbf{b}]||,$ | | , (43)
@@
As $\mathbf{P}_{\mathcal{S}}$ is diagonal, $\hat{\mathbf{P}}_{\mathcal{S}}$ can be recognized as a permutation matrix, which is decomposable into $\hat{\mathbf{P}}_{\mathcal{S}}=\prod_{j}^{d}\hat{\mathbf{P}}_{\mathcal{S}_{j}}$. Each $\hat{\mathbf{P}}_{\mathcal{S}_{j}}$ can be implemented with one multi-controlled X-gate (MCX). As MCX gates can be synthesized with $\mathcal{O}(n)$ controlled-not (CNOT) gates with a circuit depth of $\mathcal{O}(\text{log}_{2}n)$ using $\mathcal{O}(n)$ auxiliary qubits [66, 67, 68], the overall gate complexity when implementing $\hat{\mathbf{P}}_{\mathcal{S}}$ is $\mathcal{O}(dn)$ with a circuit depth of $\mathcal{O}(d\text{log}_{2}n)$. Hence, if either $d=\mathcal{O}(\text{polylog}(L))$ or $L-d=\mathcal{O}(\text{polylog}(L))$, the implementation of $\hat{\mathbf{P}}_{\mathcal{S}}$ is guaranteed to be efficient.
As future QCs might allow for implementing hardware native MCX-gates, further reductions in gate complexity are possible [69, 70, 71]. Additionally, regular patterns in $\hat{\mathbf{P}}_{\mathcal{S}}$ can be exploited to implement multiple $\hat{\mathbf{P}}_{\mathcal{S}_{j}}$ at once. Even for large $d$, this can allow for an efficient implementation of $\hat{\mathbf{P}}_{\mathcal{S}}$ in some instances, e.g., for projections to the first half of the domain.
@@
$\hat{\mathbf{O}}_{4L\times 4L}=\begin{bmatrix}\mathds{1}&-\mathds{1}&0&0\\ -\mathds{1}&\mathds{1}&0&0\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix}.$ (45)
Using the definition of |$\ket{\psi}$ detailed in equation (43), it is easy to check that
$l_{2,\mathcal{S}}^{2}=\braket{\psi}{\hat{\mathbf{O}}}{\psi}||[\mathbf{a};\mathbf{b}]||^{2}.$ (46)
The augmented observable $\hat{\mathbf{O}}$ is always - independent of the state dimension $L$ and the subspace $\mathcal{S}$ represented by the 4 Pauli matrices
$\begin{split}\hat{\mathbf{O}}_{1}&=I\otimes I\otimes I\otimes\cdots\otimes I\\ \hat{\mathbf{O}}_{2}&=I\otimes X\otimes I\otimes\cdots\otimes I\\ \hat{\mathbf{O}}_{3}&=Z\otimes I\otimes I\otimes\cdots\otimes I\\ \hat{\mathbf{O}}_{4}&=Z\otimes X\otimes I\otimes\cdots\otimes I\end{split},$ (47)
@@
$X=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},Z=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix},$ (48)
@@
$\hat{\mathbf{O}}=\sum_{j=1}^{4}\mathbf{c}_{j}\hat{\mathbf{O}}_{j}.$ (49)
@@
$l_{2,\mathcal{S}}^{2}=||[\mathbf{a};\mathbf{b}]||^{2}\sum_{j=1}^{4}\mathbf{c}_{j}\braket{\psi}{\hat{\mathbf{O}}_{j}}{\psi}.$ (50)
As proven in [72], the measurement of the expected value of a quantum state with respect to an observable that is unitary using high-fidelity amplitude amplification techniques has optimal precision scaling with a query complexity of $\mathcal{O}(\log(1/\delta)/\epsilon)$. As all Pauli observables are unitary and the number of Pauli observables that represent $\hat{\mathbf{O}}$ is independent of $L$ and $\mathcal{S}$, it follows that our algorithm has the same query complexity.
@@ A.2 Proof of Theorem 2
|$\ket{\phi}=[\mathbf{x}_{1};\mathbf{x}_{2};\dots;\mathbf{x}_{M}]/||[\mathbf{x}_{1};\mathbf{x}_{2};\dots;\mathbf{x}_{M}]||,$ (51)
with $\mathbf{x}_{i}\in\mathbb{C}^{L}$, where $M$ is the number of sub-states. We aim to estimate a squared $l_{2}$-norm, $l^{2}_{\Sigma}$, on a subspace $\mathcal{S}$ defined by the projector $\mathbf{P}_{\mathcal{S}}$, such that
$l_{\Sigma,\mathcal{S}}^{2}=||\mathbf{P}_{\mathcal{S}}\mathbf{x}_{1}+\mathbf{P}_{\mathcal{S}}\mathbf{x}_{2}+\dots+\mathbf{P}_{\mathcal{S}}\mathbf{x}_{M}||^{2}.$ (52)
It remains to generalise $\hat{\mathbf{P}}_{\mathcal{S}}$ (see equation (42)) and $\hat{\mathbf{O}}$ (see equation (49)).
We extend $\hat{\mathbf{P}}_{\mathcal{S}}$ by recognizing that it corresponds to the unitary matrix
$\hat{\mathbf{P}}_{\mathcal{S},A}=\begin{bmatrix}\mathbf{P}_{\mathcal{S}}&(\mathds{1}-\mathbf{P}_{\mathcal{S}})\\ (\mathds{1}-\mathbf{P}_{\mathcal{S}})&\mathbf{P}_{\mathcal{S}}\end{bmatrix},$ (53)
@@
$\hat{\mathbf{P}}_{\mathcal{S},\Sigma}=\hat{\mathbf{P}}_{\mathcal{S},A}\otimes I^{\otimes log_{2}M}.$ (54)
Similarly, the specific structure of $\hat{\mathbf{O}}$ lets us define the generalized observable $\hat{\mathbf{O}}_{\Sigma}$ as
$\hat{\mathbf{O}}_{\Sigma}=\frac{1}{2}(I\otimes\hat{\mathbf{O}}_{A}+Z\otimes\hat{\mathbf{O}}_{A})\otimes I^{\otimes log_{2}L},$ (55)
@@
$\hat{\mathbf{O}}_{A}=\sum_{\mathbf{A}\in\{I,X\}^{\otimes log_{2}M}}\mathbf{A}.$ (56)
In this context, equation (55) represents the half-domain measurement of |$\ket{\psi}$ through
@@
and the dimension of each sub-state $\mathbf{x}_{i}$ as $\dim\left(I^{\otimes\log_{2}L}\right)=L\times L$, while equation (56) describes the summation of the sub-states.

### v3

@@ Appendix A Estimating Loss Functions
To prove Theorem 2 we first introduce the following auxiliary result:
@@
Let $\mathbf{a},\mathbf{b}\in\mathbb{C}^{L}$. Given a projector $\mathbf{P}_{\mathcal{S}}=\text{diag}(p_{1},p_{2},\dots,p_{L})$, which projects onto a subspace $\mathcal{S}\subseteq\mathbb{C}^{L}$, where each $p_{i}\in\{0,1\}$ and $d=\sum_{i=1}^{L}p_{i}$. Let there be a state preparation oracle $\mathcal{U}$ that prepares the quantum state |$\ket{\phi}=[\mathbf{a};\mathbf{b}]/||[\mathbf{a};\mathbf{b}]||$ |. Then, if either $d=\mathcal{O}(\text{polylog}(L))$ or $L-d=\mathcal{O}(\text{polylog}(L))$, there exists a quantum algorithm that efficiently estimates the $l_{2,\mathcal{S}}$-norm between $\mathbf{a}$ and $\mathbf{b}$ on $\mathcal{S}$ as
@@ A.1 Proof of Lemma 4
By introducing |$\ket{\phi}=[\mathbf{a};\mathbf{b}]/||[\mathbf{a};\mathbf{b}]||$ |, we can write Eq. (39) as an expected value estimation
@@
To estimate $l_{2,\mathcal{S}}^{2}$ with optimal precision scaling beyond the central limit, we will introduce an auxiliary qubit to double the state space dimension. In this augmented space we can find a unitary decomposable observable $\tilde{\mathbf{O}}$, whose expectation value with respect to the augmented state |$\ket{\psi}=[\mathbf{P}_{\mathcal{S}}\mathbf{a};\mathbf{P}_{\mathcal{S}}\mathbf{b};(\mathds{1}-\mathbf{P}_{\mathcal{S}})\mathbf{a};(\mathds{1}-\mathbf{P}_{\mathcal{S}})\mathbf{b}]/||[\mathbf{a};\mathbf{b}]||$ | is equal to $l_{2,\mathcal{S}}$, such that
$l_{2,\mathcal{S}}^{2}=\braket{\psi}{\tilde{\mathbf{O}}}{\psi}||[\mathbf{a};\mathbf{b}]||^{2}.$ (42)
We first demonstrate that there exists a unitary initialization procedure for |$\ket{\psi}$ given an initialization oracle for |$\ket{\phi}$ and discuss the efficiency of its implementation. We then explicitly construct $\tilde{\mathbf{O}}$, which can be decomposed into $4$ commuting unitary observables independent of $L$ and $\mathcal{S}$. We conclude by analyzing the overall complexity of our algorithm.
@@
where $n=\log_{2}(L)$. We further define the unitary matrix $\tilde{\mathbf{P}}_{\mathcal{S}}\in\mathbb{C}^{4L\times 4L}$ constructed from $\mathbf{P}_{\mathcal{S}}$ and its complement subspace $\mathds{1}-\mathbf{P}_{\mathcal{S}}$ as
$\tilde{\mathbf{P}}_{\mathcal{S}}=\begin{bmatrix}\mathbf{P}_{\mathcal{S}}&0&(\mathds{1}-\mathbf{P}_{\mathcal{S}})&0\\ 0&\mathbf{P}_{\mathcal{S}}&0&(\mathds{1}-\mathbf{P}_{\mathcal{S}})\\ (\mathds{1}-\mathbf{P}_{\mathcal{S}})&0&\mathbf{P}_{\mathcal{S}}&0\\ 0&(\mathds{1}-\mathbf{P}_{\mathcal{S}})&0&\mathbf{P}_{\mathcal{S}}\end{bmatrix}.$ (44)
@@
|$\ket{\psi}$ $=\tilde{\mathbf{P}}_{\mathcal{S}}(I\otimes\mathbf{U})\ket{0^{n+1}}$ (45)
$=[\mathbf{P}_{\mathcal{S}}\mathbf{a};\mathbf{P}_{\mathcal{S}}\mathbf{b};(\mathds{1}-\mathbf{P}_{\mathcal{S}})\mathbf{a};(\mathds{1}-\mathbf{P}_{\mathcal{S}})\mathbf{b}]/||[\mathbf{a};\mathbf{b}]||,$
@@
As $\mathbf{P}_{\mathcal{S}}$ is diagonal, $\tilde{\mathbf{P}}_{\mathcal{S}}$ can be recognized as a permutation matrix, which is decomposable into $\tilde{\mathbf{P}}_{\mathcal{S}}=\prod_{j}^{d}\tilde{\mathbf{P}}_{\mathcal{S}_{j}}$. Each $\tilde{\mathbf{P}}_{\mathcal{S}_{j}}$ can be implemented with one multi-controlled X-gate (MCX). As MCX gates can be synthesized with $\mathcal{O}(n)$ controlled-not (CNOT) gates with a circuit depth of $\mathcal{O}(\log_{2}(n))$ using $\mathcal{O}(n)$ auxiliary qubits [70, 71, 72], the overall gate complexity when implementing $\tilde{\mathbf{P}}_{\mathcal{S}}$ is $\mathcal{O}(dn)$ with a circuit depth of $\mathcal{O}(d\log_{2}(n))$. Hence, if either $d=\mathcal{O}(\text{polylog}(L))$ or $L-d=\mathcal{O}(\text{polylog}(L))$, the implementation of $\tilde{\mathbf{P}}_{\mathcal{S}}$ is guaranteed to be efficient.
As future QC might allow for implementing hardware native MCX-gates, further reductions in gate complexity are possible [73, 74, 75]. Additionally, regular patterns in $\tilde{\mathbf{P}}_{\mathcal{S}}$ can be exploited to implement multiple $\tilde{\mathbf{P}}_{\mathcal{S}_{j}}$ at once. Even for large $d$, this can allow for an efficient implementation of $\tilde{\mathbf{P}}_{\mathcal{S}}$ in some instances, e.g., for projections to the first half of the domain.
@@
$\tilde{\mathbf{O}}_{4L\times 4L}=\begin{bmatrix}\mathds{1}&-\mathds{1}&0&0\\ -\mathds{1}&\mathds{1}&0&0\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix}.$ (47)
Using the definition of |$\ket{\psi}$, it is easy to check that
$l_{2,\mathcal{S}}^{2}=\braket{\psi}{\tilde{\mathbf{O}}}{\psi}||[\mathbf{a};\mathbf{b}]||^{2}.$ (48)
The augmented observable $\tilde{\mathbf{O}}$ is always - independent of the state dimension $L$ and the subspace $\mathcal{S}$ represented by the 4 Pauli matrices
$\begin{split}\tilde{\mathbf{O}}_{1}&=I\otimes I\otimes I\otimes\cdots\otimes I\\ \tilde{\mathbf{O}}_{2}&=I\otimes X\otimes I\otimes\cdots\otimes I\\ \tilde{\mathbf{O}}_{3}&=Z\otimes I\otimes I\otimes\cdots\otimes I\\ \tilde{\mathbf{O}}_{4}&=Z\otimes X\otimes I\otimes\cdots\otimes I\end{split},$ (49)
@@
$X=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},\quad Z=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix},$ (50)
@@
$\tilde{\mathbf{O}}=\sum_{j=1}^{4}\mathbf{c}_{j}\tilde{\mathbf{O}}_{j}.$ (51)
@@
$l_{2,\mathcal{S}}^{2}=||[\mathbf{a};\mathbf{b}]||^{2}\sum_{j=1}^{4}\mathbf{c}_{j}\braket{\psi}{\tilde{\mathbf{O}}_{j}}{\psi}.$ (52)
As proven in [76], the measurement of the expected value of a quantum state with respect to an observable that is unitary using high-fidelity amplitude amplification techniques has optimal precision scaling with a query complexity of $\mathcal{O}(\log(1/\delta)/\epsilon)$. As all Pauli observables are unitary and the number of Pauli observables that represent $\tilde{\mathbf{O}}$ is independent of $L$ and $\mathcal{S}$, it follows that our algorithm has the same query complexity.
@@ A.2 Proof of Theorem 2
|$\ket{\phi}=[\mathbf{w}_{1};\mathbf{w}_{2};\dots;\mathbf{w}_{M}]/||[\mathbf{w}_{1};\mathbf{w}_{2};\dots;\mathbf{w}_{M}]||,$ (53)
with $\mathbf{w}_{i}\in\mathbb{C}^{L}$, where $M$ is the number of sub-states. We aim to estimate a squared $l_{2}$-norm, $l^{2}_{\Sigma}$, on a subspace $\mathcal{S}$ defined by the projector $\mathbf{P}_{\mathcal{S}}$, such that
$l_{\Sigma,\mathcal{S}}^{2}=||\mathbf{P}_{\mathcal{S}}\mathbf{w}_{1}+\mathbf{P}_{\mathcal{S}}\mathbf{w}_{2}+\dots+\mathbf{P}_{\mathcal{S}}\mathbf{w}_{M}||^{2}.$ (54)
It remains to generalise $\tilde{\mathbf{P}}_{\mathcal{S}}$ (see Eq. (44)) and $\tilde{\mathbf{O}}$ (see Eq. (51)).
We extend $\tilde{\mathbf{P}}_{\mathcal{S}}$ by recognizing that it corresponds to the unitary matrix
$\tilde{\mathbf{P}}_{\mathcal{S},A}=\begin{bmatrix}\mathbf{P}_{\mathcal{S}}&(\mathds{1}-\mathbf{P}_{\mathcal{S}})\\ (\mathds{1}-\mathbf{P}_{\mathcal{S}})&\mathbf{P}_{\mathcal{S}}\end{bmatrix},$ (55)
@@
$\tilde{\mathbf{P}}_{\mathcal{S},\Sigma}=\tilde{\mathbf{P}}_{\mathcal{S},A}\otimes I^{\otimes\log_{2}(M)}.$ (56)
Similarly, the specific structure of $\tilde{\mathbf{O}}$ lets us define the generalized observable $\tilde{\mathbf{O}}_{\Sigma}$ as
$\tilde{\mathbf{O}}_{\Sigma}=\frac{1}{2}(I\otimes\tilde{\mathbf{O}}_{A}+Z\otimes\tilde{\mathbf{O}}_{A})\otimes I^{\otimes\log_{2}(L)},$ (57)
@@
$\tilde{\mathbf{O}}_{A}=\sum_{\mathbf{A}\in\{I,X\}^{\otimes\log_{2}(M)}}\mathbf{A}.$ (58)
In this context, Eq. (57) represents the half-domain measurement of |$\ket{\psi}$ through
@@
and the dimension of each sub-state $\mathbf{w}_{i}$ as $\dim\left(I^{\otimes\log_{2}(L)}\right)=L\times L$, while Eq. (58) describes the summation of the sub-states.